### v3

@@ 1 Introduction
Transmission lines are among the most spatially distributed and environmentally exposed assets in modern power grids, continuously subjected to weathering, cyclic mechanical loading, and progressive material degradation such as corrosion and insulation aging (Zhang et al., 2024b; Luo et al., 2023). Defects arising from these processes—including insulator contamination, conductor corrosion, and fitting damage—as well as external hazards such as bird nesting, can lead to flashovers, line tripping, and unplanned outages, thereby increasing maintenance costs and compromising grid reliability (Bao et al., 2022; Cheng and Liu, 2024). Timely identification of such defects is therefore critical for prioritizing maintenance scheduling, reducing unnecessary patrols, and mitigating safety risks (Yang et al., 2023). Because transmission networks often extend over thousands of kilometers, this task calls for automated visual inspection methods that can operate at high frequency and low marginal cost—requirements that are difficult to meet through manual or semi-automated workflows owing to limited workforce availability, access constraints in remote corridors, and the sheer volume of infrastructure to be covered.
@@
Conventional inspection relies on human operators traversing transmission corridors on foot or via helicopter, or deploying robotic platforms such as line-crawling robots along predefined routes (Pouliot and Montambault, 2012; Xu et al., 2022a). These approaches are hazardous, labor-intensive, and costly to sustain at the scale of a national grid. The widespread adoption of unmanned aerial vehicles (UAVs) equipped with high-resolution cameras has substantially improved data acquisition, enabling efficient imagery collection over long corridors (Xu et al., 2022b; da Silva et al., 2020) and motivating a shift toward learning-based defect detection. The imagery obtained in this manner, however, poses detection challenges that differ markedly from those encountered in general-purpose benchmarks such as COCO or VisDrone. Transmission-line scenes are dominated by repetitive structural patterns and geometrically elongated components whose dense edge responses and strong background textures severely compress the effective pixel area available for defect discrimination. Defects typically occupy only a small fraction of the image, and this difficulty is amplified by oblique viewing angles, varying shooting distances, and persistent background clutter from vegetation, sky, and metallic infrastructure. On the TL-RGBD benchmark adopted in this study, approximately 94.5% of annotated defect instances fall below the COCO small-object area threshold of $32\times 32$ pixels at the $640\times 640$ training input resolution, indicating that detection methods not explicitly designed for small objects tend to underperform in this domain.
@@
Deep learning has driven substantial progress in transmission-line defect detection by exploiting the growing volume of UAV-acquired data. Existing approaches rely almost exclusively on RGB imagery as the sole input modality. Representative works span YOLO variants with attention-augmented backbones (Wang et al., 2024), Faster R-CNN enhanced with aerial-image augmentation strategies (Cao et al., 2023), lightweight architectures with multiscale feature fusion (Ma et al., 2025; Wu et al., 2025), and self-supervised vision transformer pipelines (Zhang et al., 2024a). A shared limitation of these methods is their dependence on photometric appearance—texture, color, and shading—as the primary discriminative cue. This dependence becomes a major bottleneck when defects exhibit low chromatic contrast against cluttered backgrounds, when damage patterns are geometrically ambiguous, or when targets are partially occluded by vegetation (He et al., 2025; Chen et al., 2025). Because RGB images encode geometry only implicitly through shading and perspective, rather than providing explicit three-dimensional surface cues, separating subtle structural deformations from benign surface variations is particularly challenging under such conditions (Li et al., 2025; Gupta et al., 2014). An additional accuracy–efficiency tension constrains practical deployment: high-capacity models impose computational costs that exceed the processing budget of typical UAV-mounted hardware at the frame rates required for real-time inspection, whereas lightweight alternatives tend to sacrifice detection sensitivity on the small targets that dominate real-world transmission-line imagery (Tan et al., 2020; Chen et al., 2024).
@@
Integrating depth information with RGB appearance offers a promising path toward more discriminative representations. Depth maps encode relative surface geometry and spatial extent independently of illumination and texture, thereby providing complementary evidence where photometric cues alone are insufficient—for instance, by revealing surface discontinuities at damaged fittings or volumetric protrusions caused by bird nests. However, fusing heterogeneous modalities is not straightforward. RGB and depth features originate from fundamentally different sensing processes and exhibit distinct noise characteristics and distributional properties: depth maps may contain holes, quantization artifacts, and edge bleeding, while RGB images are susceptible to specular highlights and illumination variation. Direct concatenation or element-wise combination risks propagating these modality-specific artifacts into the joint representation, ultimately degrading rather than improving detection accuracy (Zhao et al., 2024). How to align and fuse such heterogeneous representations—particularly under the combined pressures of modality-specific noise and the extreme small-object prevalence that characterizes transmission-line inspection—remains insufficiently addressed for this domain.
@@
To address these challenges, this paper proposes CMAFNet, a Cross-Modal Alignment and Fusion Network built upon a purify-then-fuse paradigm for RGB-D feature integration. The design is motivated by the observation that multimodal fusion operating on noisy, statistically misaligned representations tends to inherit and amplify their deficiencies rather than compensate for them. CMAFNet therefore interposes a dedicated purification stage before fusion: within each modality branch, a Semantic Recomposition Module (SRM) reconstructs input features through a structured bottleneck that suppresses modality-specific artifacts—such as depth holes and edge bleeding, or RGB specular noise—and reduces distributional mismatch between the two feature streams prior to integration. Fusion then operates on these purified representations, and a Contextual Semantic Integration Framework (CSIF) captures long-range structural dependencies that convolutional operators model only at disproportionate cost—a capability important for reasoning about spatial relationships among small defects embedded in densely repetitive structural patterns. To mitigate the detail erosion commonly associated with global attention, CSIF adopts a partial-channel design that balances global context modeling with the fine-grained spatial discrimination required for accurate small-object localization. The resulting architecture is instantiated as a scalable model family, ranging from compact configurations suitable for real-time UAV inference to full-scale variants for server-side post-processing. Experiments on the TL-RGBD benchmark show that CMAFNet consistently outperforms compared methods across evaluation metrics and model scales; for example, the full-scale variant achieves a gain of +2.8 mAP50 and +3.6 mAPS over the strongest baseline under an identical training protocol, while the lightweight variant maintains competitive accuracy at real-time throughput.
The main contributions of this paper are as follows:
  1. 1.
    We introduce the purify-then-fuse paradigm for RGB-D feature integration in small-object-dominated detection tasks. This paradigm is instantiated as CMAFNet, a scalable architecture family spanning lightweight to full-scale configurations, which achieves consistent improvements across all model scales on TL-RGBD. Systematic fusion-strategy ablations confirm that purification before fusion yields measurable gains over direct-fusion alternatives.
  2. 2.
    We propose the Semantic Recomposition Module (SRM), which suppresses modality-specific noise—including depth quantization artifacts and RGB illumination variations—and reduces cross-modal distributional mismatch through a structured reconstruction bottleneck. Applied within each modality branch and at fused feature levels, SRM brings heterogeneous RGB and depth representations into more compatible statistical properties, as validated by modality ablations and noise-sensitivity analyses presented in Section 3.
  3. 3.
    We design the Contextual Semantic Integration Framework (CSIF), a partial-channel global attention mechanism that captures long-range structural dependencies while preserving the spatial detail essential for small-object detection. By restricting global attention to a subset of feature channels and retaining the remainder for local discrimination, CSIF addresses the tension between contextual coverage and spatial precision that limits conventional full-channel attention, as demonstrated through component-level ablation.
@@ 2.1 Overview of Network Architecture
Although integrating depth geometry with RGB appearance promises to strengthen defect detection, realizing this potential hinges on how the two modalities interact during feature extraction and fusion. Naively combining them—whether by early concatenation or late decision fusion—yields limited gains, because RGB and depth signals differ substantially in statistical properties and noise characteristics (Zhao et al., 2024). Early concatenation forces the network to reconcile heterogeneous distributions from the first layer onward, while late fusion defers cross-modal interaction to the prediction stage and affords little opportunity for mutual enhancement. A more principled strategy is to fuse modality-specific features at intermediate representation levels, where sufficient semantic abstraction exists to tolerate noise yet spatial resolution remains adequate for small-object localization. Existing approaches that operate at this level, however, typically employ static operations such as element-wise summation or channel concatenation followed by convolution. Such fixed fusion schemes cannot adapt to the dynamic variations inherent in aerial inspection: when RGB features are corrupted by illumination artifacts or depth features contain boundary discontinuities and quantization noise, static fusion propagates these modality-specific corruptions into the shared representation, degrading detection performance rather than enhancing it.
These considerations motivate a fusion paradigm in which each modality stream first undergoes explicit noise suppression and distribution normalization before cross-modal integration, and in which the fused representation is subsequently refined through context-aware mechanisms that extend beyond the local receptive field of convolutional operations. The proposed Cross-Modal Alignment and Fusion Network (CMAFNet) instantiates this paradigm by extending the YOLO 11(Jocher et al., 2023) single-stream detector into a dual-branch architecture for RGB-D input, retaining all standard components—C3k2, SPPF, the FPN-PAN neck, and the anchor-free decoupled detection head—without modification. Three design elements distinguish CMAFNet from the baseline: (i) a Semantic Recomposition Module (SRM) that performs bottleneck-based feature purification with position-wise distribution normalization, (ii) a Contextual Semantic Integration Framework (CSIF) that introduces attention-driven global context modeling at the deepest fusion level, and (iii) a selective fusion strategy that restricts cross-modal integration to $P_{4}$ and $P_{5}$ while routing $P_{3}$ features exclusively from the RGB branch. Fig. 1 illustrates the complete topology.
@@
The data flow comprises three stages. In the first, the dual-branch backbone encodes RGB and depth inputs independently through to $P_{5}$, with SRM inserted at $P_{3}$ and $P_{4}$ within each branch to suppress modality-specific noise and narrow the inter-modal distribution gap. In the second, cross-modal fusion at $P_{4}$ and $P_{5}$ combines the purified branch features through channel concatenation, C3k2 re-encoding, and a second SRM pass; at $P_{5}$, CSIF is additionally interposed between C3k2 and the terminal SRM to model global context. In the third, the detection head receives the fused $P_{4}$$P_{5}$ representations together with the RGB-only $P_{3}$ features through the standard FPN-PAN bidirectional pyramid for multi-scale prediction.
@@ 2.1.1 Backbone: Dual Parallel Branches
The backbone comprises two encoding branches with identical topology but entirely separate parameters, one for RGB and one for depth. All convolutional layers, C3k2 blocks, and the terminal SPPF module adopt the YOLO 11(Jocher et al., 2023) default implementation without modification.
The RGB branch receives 3-channel images. The depth branch receives single-channel depth maps.Depth maps in the TL-RGBD dataset are spatially registered with their RGB counterparts via the factory calibration of the onboard RGB-D sensor (Intel RealSense D435i), and no additional geometric alignment is performed during preprocessing. Missing depth values and edge artifacts inherent to the sensor are not explicitly inpainted; robustness to these corruptions is instead delegated to SRM purification and the selective fusion strategy described in Section 2.1.2.
For the nano-scale configuration (width_multiple =0.25=0.25, depth_multiple =0.50=0.50, input $640\times 640$), the feature hierarchy spans five pyramid levels: $P_{1}$ ($320{\times}320$, C=16C{=}16), $P_{2}$ ($160{\times}160$, C=32C{=}32), $P_{3}$ ($80{\times}80$, C=64C{=}64), $P_{4}$ ($40{\times}40$, C=128C{=}128), and $P_{5}$ ($20{\times}20$, C=256C{=}256), with strides of 2, 4, 8, 16, and 32 respectively. The two branches produce identical feature dimensions at every level; the sole asymmetry is that the depth branch’s $P_{3}$ output is not propagated to the detection head.
Within each branch, an SRM instance (K=128K{=}128, $\alpha{=}0.8$; see Section 2.2) is inserted immediately after the C3k2 block at both $P_{3}$ and $P_{4}$. These scales strike a favorable balance: semantic content is rich enough for effective refinement, yet spatial resolution is still fine enough to preserve small-defect detail. Applying SRM before fusion narrows the inter-modal distribution gap at the source, providing cleaner inputs for the subsequent concatenation-based integration.
The rationale for parallel branches, rather than early fusion, lies in the distinct noise profiles of the two modalities. RGB imagery exhibits illumination variation and specular reflections; depth maps suffer from edge bleeding, missing values at reflective surfaces, and quantization noise. Merging these signals before individual processing would entangle their noise distributions and complicate subsequent alignment. Independent encoding through $P_{5}$ lets each branch develop modality-appropriate representations, deferring cross-modal integration to $P_{4}$$P_{5}$ where greater semantic abstraction makes noise-induced misalignment less problematic.
@@ 2.1.2 Fusion Layers: Selective Cross-Modal Integration at $P_{4}$ and $P_{5}$
Two observations motivate excluding depth from $P_{3}$. First, depth features at this resolution carry the highest sensor noise and spatial misalignment; injecting them into the feature map most critical for small-defect localization would introduce more interference than useful geometric cues. Second, depth primarily contributes structural context—surface discontinuities and inter-component spatial layout—that is adequately captured at $P_{4}$ and $P_{5}$, where feature granularity better matches the sensor’s effective resolution.

### v4

@@ 1 Introduction
Transmission lines are among the most spatially distributed and environmentally exposed assets in modern power grids, continuously subjected to weathering, cyclic mechanical loading, and progressive material degradation such as corrosion and insulation aging (Zhang et al., 2024b; Luo et al., 2023). Defects arising from these processes, including insulator contamination, conductor corrosion, and fitting damage, as well as external hazards such as bird nesting, can lead to flashovers, line tripping, and unplanned outages, increasing maintenance costs and compromising grid reliability (Bao et al., 2022; Cheng and Liu, 2024). Timely identification of such defects is therefore critical for prioritizing maintenance scheduling, reducing unnecessary patrols, and mitigating safety risks (Yang et al., 2023). Because transmission networks often extend over thousands of kilometers, this task calls for automated visual inspection methods that can operate at high frequency and low marginal cost, requirements that are difficult to meet through manual or semi-automated workflows given limited workforce availability, access constraints in remote corridors, and the sheer volume of infrastructure to be covered.
@@
Conventional inspection relies on human operators traversing transmission corridors on foot or by helicopter, or deploying robotic platforms such as line-crawling robots along predefined routes (Pouliot and Montambault, 2012; Xu et al., 2022a). These approaches are hazardous, labor-intensive, and costly to sustain at national-grid scale. The widespread adoption of unmanned aerial vehicles (UAVs) equipped with high-resolution cameras has substantially improved data acquisition, enabling efficient imagery collection over long corridors (Xu et al., 2022b; da Silva et al., 2020) and motivating a shift toward learning-based defect detection. The imagery obtained in this way, however, poses detection challenges that differ markedly from those in general-purpose benchmarks such as COCO or VisDrone. Transmission-line scenes are dominated by repetitive structural patterns and geometrically elongated components whose dense edge responses and strong background textures severely compress the effective pixel area available for defect discrimination. Defects typically occupy only a small fraction of the image, and this difficulty is compounded by oblique viewing angles, varying shooting distances, and persistent background clutter from vegetation, sky, and metallic infrastructure. On the TL-RGBD benchmark adopted in this study, approximately 94.5% of annotated defect instances fall below the COCO small-object area threshold of $32\times 32$ pixels at the $640\times 640$ training input resolution, confirming that detection methods not explicitly designed for small objects tend to underperform in this domain.
@@
Deep learning has driven substantial progress in transmission-line defect detection by exploiting the growing volume of UAV-acquired data. Existing approaches rely almost exclusively on RGB imagery as the sole input modality. Representative works span YOLO variants with attention-augmented backbones (Wang et al., 2024), Faster R-CNN enhanced with aerial-image augmentation strategies (Cao et al., 2023), lightweight architectures with multiscale feature fusion (Ma et al., 2025; Wu et al., 2025), and self-supervised vision transformer pipelines (Zhang et al., 2024a). A shared limitation of these methods is their dependence on photometric appearance, namely texture, color, and shading, as the primary discriminative cue. This dependence becomes a bottleneck when defects exhibit low chromatic contrast against cluttered backgrounds, when damage patterns are geometrically ambiguous, or when targets are partially occluded by vegetation (He et al., 2025; Chen et al., 2025). Because RGB images encode geometry only implicitly through shading and perspective rather than providing explicit three-dimensional surface cues, separating subtle structural deformations from benign surface variations remains difficult under such conditions (Li et al., 2025; Gupta et al., 2014). A further accuracy-efficiency tension constrains practical deployment: high-capacity models impose computational costs that exceed the processing budget of typical UAV-mounted hardware at the frame rates required for real-time inspection, while lightweight alternatives tend to sacrifice detection sensitivity on the small targets that dominate real-world transmission-line imagery (Tan et al., 2020; Chen et al., 2024).
@@
Adding depth information to RGB appearance offers a promising path toward more discriminative representations. Depth maps encode relative surface geometry and spatial extent independently of illumination and texture, providing complementary evidence where photometric cues alone fall short, for instance by revealing surface discontinuities at damaged fittings or volumetric protrusions caused by bird nests. Fusing heterogeneous modalities, however, is not straightforward. RGB and depth features originate from fundamentally different sensing processes and exhibit distinct noise characteristics and distributional properties: depth maps may contain holes, quantization artifacts, and edge bleeding, while RGB images are susceptible to specular highlights and illumination variation. Direct concatenation or element-wise combination risks propagating these modality-specific artifacts into the joint representation, ultimately degrading rather than improving detection accuracy (Zhao et al., 2024). How to align and fuse such heterogeneous representations, particularly under the combined pressures of modality-specific noise and the extreme small-object prevalence that characterizes transmission-line inspection, remains insufficiently addressed for this domain.
@@
To address these challenges, this paper proposes CMAFNet, a Cross-Modal Alignment and Fusion Network built on a purify-then-fuse paradigm for RGB-D feature integration. The design is motivated by the observation that multimodal fusion operating on noisy, statistically misaligned representations tends to inherit and amplify their deficiencies rather than compensate for them. CMAFNet therefore interposes a dedicated purification stage before fusion: within each modality branch, a Semantic Recomposition Module (SRM) reconstructs input features through a structured bottleneck that suppresses modality-specific artifacts, such as depth holes and edge bleeding or RGB specular noise, and reduces distributional mismatch between the two feature streams prior to integration. Fusion then operates on these purified representations, and a Contextual Semantic Integration Framework (CSIF) captures long-range structural dependencies that convolutional operators model only at disproportionate cost, a capability important for reasoning about spatial relationships among small defects embedded in densely repetitive structural patterns. To mitigate the detail erosion commonly associated with global attention, CSIF adopts a partial-channel design that balances global context modeling with the fine-grained spatial discrimination required for accurate small-object localization. The resulting architecture is instantiated as a scalable model family, ranging from compact configurations suitable for real-time UAV inference to full-scale variants for server-side post-processing. Experiments on the TL-RGBD benchmark show that CMAFNet consistently outperforms compared methods across evaluation metrics and model scales; for example, the full-scale variant achieves a gain of +2.8 mAP50 and +3.6 mAPS over the strongest baseline under an identical training protocol, while the lightweight variant maintains competitive accuracy at real-time throughput.
The core contributions of this work are threefold. We introduce the purify-then-fuse paradigm for RGB-D feature integration in detection tasks dominated by small objects, instantiated as CMAFNet, a scalable architecture family spanning lightweight to full-scale configurations that achieves consistent improvements across all model scales on TL-RGBD; systematic fusion-strategy ablations confirm that purification before fusion yields measurable gains over direct-fusion alternatives. We propose the Semantic Recomposition Module (SRM), which suppresses modality-specific noise, including depth quantization artifacts and RGB illumination variations, and reduces cross-modal distributional mismatch through a structured reconstruction bottleneck; applied within each modality branch and at fused feature levels, SRM brings heterogeneous RGB and depth representations into more compatible statistical properties, as validated by modality ablations and noise-sensitivity analyses presented in Section 3. We also design the Contextual Semantic Integration Framework (CSIF), a partial-channel global attention mechanism that captures long-range structural dependencies while preserving the spatial detail essential for small-object detection; by restricting global attention to a subset of feature channels and retaining the remainder for local discrimination, CSIF addresses the tension between contextual coverage and spatial precision that limits conventional full-channel attention, as shown through component-level ablation.
@@ 2.1 Overview of Network Architecture
Although integrating depth geometry with RGB appearance holds promise for strengthening defect detection, the effectiveness of such integration depends critically on how the two modalities interact during feature extraction and fusion. Naive combination strategies—whether through early concatenation or late decision fusion—yield limited improvements, as RGB and depth signals differ substantially in statistical properties and noise characteristics (Zhao et al., 2024). Early concatenation forces the network to reconcile heterogeneous distributions from the first layer onward, while late fusion defers cross-modal interaction to the prediction stage and provides little opportunity for mutual enhancement. A more principled strategy is to fuse modality-specific features at intermediate representation levels, where sufficient semantic abstraction exists to tolerate noise while spatial resolution remains adequate for small-object localization. Existing approaches operating at this level, however, typically employ static operations such as element-wise summation or channel concatenation followed by convolution. Such fixed fusion schemes cannot adapt to the dynamic variations inherent in aerial inspection: when RGB features are corrupted by illumination artifacts or depth features contain boundary discontinuities and quantization noise, static fusion propagates these modality-specific corruptions into the shared representation, ultimately degrading rather than enhancing detection performance.
These considerations motivate a fusion paradigm in which each modality stream first undergoes explicit noise suppression and distribution normalization prior to cross-modal integration, and in which the fused representation is subsequently refined through context-aware mechanisms that extend beyond the local receptive field of standard convolutions. The proposed Cross-Modal Alignment and Fusion Network (CMAFNet) instantiates this paradigm by extending the YOLO 11 (Jocher et al., 2023) single-stream detector into a dual-branch architecture for RGB-D input, while retaining all standard components—C3k2, SPPF, the FPN-PAN neck, and the anchor-free decoupled detection head—without modification. Three design elements distinguish CMAFNet from the baseline: (i) a Semantic Recomposition Module (SRM) that performs bottleneck-based feature purification with position-wise distribution normalization, (ii) a Contextual Semantic Integration Framework (CSIF) that introduces attention-driven global context modeling at the deepest fusion level, and (iii) a selective fusion strategy that restricts cross-modal integration to $P_{4}$ and $P_{5}$ while routing $P_{3}$ features exclusively from the RGB branch. Fig. 1 illustrates the complete network topology.
@@
The data flow comprises three stages. In the first stage, the dual-branch backbone encodes RGB and depth inputs independently through to $P_{5}$, with SRM inserted at $P_{3}$ and $P_{4}$ within each branch to suppress modality-specific noise and narrow the inter-modal distribution gap. In the second stage, cross-modal fusion at $P_{4}$ and $P_{5}$ combines the purified branch features through channel concatenation, C3k2 re-encoding, and a second SRM pass; at $P_{5}$, CSIF is additionally interposed between C3k2 and the terminal SRM to capture global context. In the third stage, the detection head receives the fused $P_{4}$$P_{5}$ representations together with the RGB-only $P_{3}$ features through the standard FPN-PAN bidirectional pyramid for multi-scale prediction.
@@ 2.1.1 Backbone: Dual Parallel Branches
The backbone comprises two encoding branches with identical topology but entirely separate parameters, one processing RGB input and the other processing depth input. All convolutional layers, C3k2 blocks, and the terminal SPPF module adopt the default YOLO 11 (Jocher et al., 2023) implementation without modification.
The RGB branch receives 3-channel images, while the depth branch receives single-channel depth maps. Depth maps in the TL-RGBD dataset are spatially registered with their RGB counterparts through the factory calibration of the onboard RGB-D sensor (Intel RealSense D435i), and no additional geometric alignment is performed during preprocessing. Missing depth values and edge artifacts inherent to the sensor are not explicitly inpainted; robustness to these corruptions is instead delegated to the SRM purification and the selective fusion strategy described in Section 2.1.2.
For the nano-scale configuration (width_multiple =0.25=0.25, depth_multiple =0.50=0.50, input $640\times 640$), the feature hierarchy spans five pyramid levels: $P_{1}$ ($320{\times}320$, C=16C{=}16), $P_{2}$ ($160{\times}160$, C=32C{=}32), $P_{3}$ ($80{\times}80$, C=64C{=}64), $P_{4}$ ($40{\times}40$, C=128C{=}128), and $P_{5}$ ($20{\times}20$, C=256C{=}256), with strides of 2, 4, 8, 16, and 32, respectively. The two branches produce identical feature dimensions at every level; the sole asymmetry is that the output of $P_{3}$ in the depth branch is not propagated to the detection head.
Within each branch, an SRM instance (K=128K{=}128, $\alpha{=}0.8$; see Section 2.2) is inserted immediately after the C3k2 block at both $P_{3}$ and $P_{4}$. These scales strike a favorable balance: semantic content is sufficiently rich for effective refinement, yet spatial resolution remains fine enough to preserve small-defect detail. Applying SRM before fusion narrows the inter-modal distribution gap at the source, thereby providing cleaner inputs for the subsequent concatenation-based integration.
The rationale for adopting parallel branches, rather than early fusion, lies in the distinct noise profiles of the two modalities. RGB imagery exhibits illumination variation and specular reflections, whereas depth maps suffer from edge bleeding, missing values at reflective surfaces, and quantization noise. Merging these signals before individual processing would entangle their noise distributions and complicate subsequent alignment. Independent encoding through $P_{5}$ allows each branch to develop modality-appropriate representations, deferring cross-modal integration to $P_{4}$ and $P_{5}$, where greater semantic abstraction renders noise-induced misalignment less problematic.
@@ 2.1.2 Fusion Layers: Selective Cross-Modal Integration at $P_{4}$ and $P_{5}$
Two observations motivate the exclusion of depth from $P_{3}$. First, depth features at this resolution carry the highest level of sensor noise and spatial misalignment; injecting them into the feature map most critical for small-defect localization would introduce more interference than useful geometric cues. Second, the primary contribution of depth—surface discontinuities and inter-component spatial layout—is adequately captured at $P_{4}$ and $P_{5}$, where feature granularity better matches the effective resolution of the sensor.